# Neuro-Symbolic Generation of Explanations for Robot Policies with Weighted Signal Temporal Logic

### v1

@@ V Conclusion
We introduced a neuro-symbolic framework, TLNet, for generating weighted Signal Temporal Logic (wSTL) explanations of robot policies, addressing limitations of trajectory classification-based methods. Our approach directly explains policies while refining explanations through predicate filtering, loss regularization, and weight pruning. To assess quality beyond classification metrics, we proposed human-interpretability metrics—conciseness, consistency, and strictness. Experiments in three simulated robotics environments showed superior explanation explanability with high accuracy, with an ablation study confirming our simplification process’s effectiveness. By bridging neural policy learning and formal logic-based explanations, our work advances transparency in robotics, with future directions including more complex wSTL structures, human-in-the-loop refinement, and real-world applications.

### v2

@@ V Conclusion
We introduce a neuro-symbolic framework, based on the TLNet architecture, for generating wSTL explanations of robot policies. Our approach generates easy-to-understand explanations through predicate filtering, loss regularization, and weight pruning. To assess explanation quality beyond classification metrics, we also propose explainability metrics—conciseness, consistency, and strictness. Experiments in three simulated robotic environments show that our method produces explanations that better capture these metrics while retaining high classification accuracy. By bridging neural network learning and formal logic-based explanations, our work advances explainability in complex robotic applications, with future directions including more expressive wSTL structures, human-in-the-loop refinement, and explainability-by-design policy optimization.